# High-order persistence of resonant caustics in perturbed circular billiards

### v1

@@ Appendix B Computation of some asymptotic expansions
(17) ~$\tilde{\mathcal{H}}_{k}=\sum_{i=1}^{k-1}\sum_{j=1}^{k-i}\left(\sum_{|{\boldsymbol{\alpha}}|=j,\|{\boldsymbol{\alpha}}\|=k-i}\frac{{\boldsymbol{\psi}}^{\boldsymbol{\alpha}}}{{\boldsymbol{\alpha}}!}\right)h_{i}^{(j)},\quad\forall k\geq 1.$
@@
~$\tilde{\mathcal{R}}_{k}$ $=c\mathcal{H}_{k}[\![h_{\leq k},\psi_{<k}]\!]+\tilde{\mathcal{C}}_{k}[\![\theta_{<k}]\!]+\textstyle\sum_{l=1}^{k-1}\mathcal{H}_{l}[\![h_{\leq l},\psi_{<l}]\!]\mathcal{C}_{k-l}[\![\theta_{\leq k-l}]\!].$
@@ Appendix C Proofs of Theorem 9 and Lemma 10
    ~$\tilde{\mathcal{R}}_{k},\mathcal{Q}_{k}\in T_{nk}[t]$ —so condition (11) holds because $q>nm$ and $\bar{\mathcal{Q}}_{k}=0$—;
@@
The base case $m=1$ is trivial. Namely, $\mathcal{H}_{1}=h_{1}\in T_{n}[t]$ by condition (3), ~$\tilde{\mathcal{R}}_{1}=ch_{1}\in T_{n}[t]$, $\mathcal{Q}_{1}=sh^{\prime}_{1}\in T_{n}[t]$, $\theta_{1}\in T_{n}[t]$ is the unique solution of problem (15), $\varphi_{1}\in T_{n}[t]$ is the unique solution of equation $\delta\{\varphi_{1}\}=2\theta_{1}$ such that $\mu\{\varphi_{1}\}=0$, $\psi_{1}=\varphi_{1}+\theta_{1}\in T_{n}[t]$, $\mathcal{S}_{1}=c\theta_{1}\in T_{n}[t]$ and $\mathcal{C}_{1}=-s\theta_{1}\in T_{n}[t]$.
@@
If condition (3) holds, $\Gamma_{\epsilon}$ is anti-centrally symmetric, $q\neq m\mod 2$, $m\geq 2$ and $q>n(m-1)$, then we can compute all $\mathop{\rm O}\nolimits(\epsilon^{m})$-corrections too.
@@
We already know from Lemma 17 that we can compute the $\mathop{\rm O}\nolimits(\epsilon^{m-1})$-corrections, since $q>n(m-1)$. Therefore, we only need to check that the last persistence condition

### v2

@@ Appendix B Computation of some asymptotic expansions
(17) $\tilde{\mathcal{H}}_{k}=\sum_{i=1}^{k-1}\sum_{j=1}^{k-i}\left(\sum_{|{\boldsymbol{\alpha}}|=j,\|{\boldsymbol{\alpha}}\|=k-i}\frac{{\boldsymbol{\psi}}^{\boldsymbol{\alpha}}}{{\boldsymbol{\alpha}}!}\right)h_{i}^{(j)},\quad\forall k\geq 1.$
@@
$\tilde{\mathcal{R}}_{k}$ $=c\mathcal{H}_{k}[\![h_{\leq k},\psi_{<k}]\!]+\tilde{\mathcal{C}}_{k}[\![\theta_{<k}]\!]+\textstyle\sum_{l=1}^{k-1}\mathcal{H}_{l}[\![h_{\leq l},\psi_{<l}]\!]\mathcal{C}_{k-l}[\![\theta_{\leq k-l}]\!].$
@@ Appendix C Proofs of Theorem 9 and Lemma 10
    $\tilde{\mathcal{R}}_{k},\mathcal{Q}_{k}\in T_{nk}[t]$ —so condition (11) holds because $q>nm$ and $\bar{\mathcal{Q}}_{k}=0$—;
@@
The base case $m=1$ is trivial. Namely, $\mathcal{H}_{1}=h_{1}\in T_{n}[t]$ by condition (3), $\tilde{\mathcal{R}}_{1}=ch_{1}\in T_{n}[t]$, $\mathcal{Q}_{1}=sh^{\prime}_{1}\in T_{n}[t]$, $\theta_{1}\in T_{n}[t]$ is the unique solution of problem (15), $\varphi_{1}\in T_{n}[t]$ is the unique solution of equation $\delta\{\varphi_{1}\}=2\theta_{1}$ such that $\mu\{\varphi_{1}\}=0$, $\psi_{1}=\varphi_{1}+\theta_{1}\in T_{n}[t]$, $\mathcal{S}_{1}=c\theta_{1}\in T_{n}[t]$ and $\mathcal{C}_{1}=-s\theta_{1}\in T_{n}[t]$.
@@
If condition (3) holds, $\Gamma_{\epsilon}$ is anti-centrally symmetric, $q\neq m\mod 2$, $m\geq 2$ and q>n(m1)q>n(m-1), then we can compute all $\mathop{\rm O}\nolimits(\epsilon^{m})$-corrections too.
@@
We already know from Lemma 17 that we can compute the $\mathop{\rm O}\nolimits(\epsilon^{m-1})$-corrections, since q>n(m1)q>n(m-1). Therefore, we only need to check that the last persistence condition